OpenStudy (anonymous):

how can you determine whether a function is even or not? e.g x^4 or x^3+1. Thanks

OpenStudy (anonymous):

even fuctions if f(x) = f(-x) odd function if f(-x) = -f(x)

OpenStudy (anonymous):

so for x^4 : 2^4 = 16 and -2^4 = 16 so x^4 is even and for x^3 + 1: 2^3 + 1 = 9 and -2^3 + 1 = -7 not even now -(2^3 + 1) = -9 and -2^3 + 1 = -7 so neither requirement is made so x^3 + 1: is neither even nor odd